• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Ramesh

    Log entry time 19:40:20 on June 2,2005

    Entry number 144908

    Followups:

    keyword=holog entries made from adaql8 did not go to the web

    I had made two halog entries from the adaql8 (at the terminal for shift leader), they did not go to the web. They were about the left spectrometer magnet-reboot. Now the web-interface-reboot seems to work fine. Left hrs magnets are stable now. I cycled left Q2 and Q3. I also ramped up all left magnets at the P0 setting 3.643 GeV. We are waiting for beam.